# Chlorophyll-a Mapping and Prediction in the Mar Menor Lagoon Using C2RCC-Processed Sentinel 2 Imagery

### v1

@@ Abstract
The Mar Menor, Europe’s largest hypersaline coastal lagoon, located in southeastern Spain, has undergone severe eutrophication crises, with devastating impacts on biodiversity and water quality. Monitoring chlorophyll-a (Chl-a), a proxy for phytoplankton biomass, is essential to anticipate harmful algal blooms and guide mitigation. Traditional in situ measurements, while precise, are spatially and temporally limited. Satellite-based approaches provide a more comprehensive view, enabling scalable, long-term, and transferable monitoring. This study aims to overcome limitations of chlorophyll monitoring, often restricted to surface estimates or limited temporal coverage, by developing a reliable methodology to predict and map Chl-a concentrations across the entire water column of the Mar Menor. Specifically, the work integrates Sentinel 2 imagery with buoy-based ground truth to create models capable of high-resolution, depth-specific monitoring, enhancing early-warning capabilities for eutrophication. Nearly a decade of Sentinel 2 images was atmospherically corrected using C2RCC processors. Buoy data were aggregated by depth (0–1 m, 1–2 m, 2–3 m, 3–4 m). Multiple machine learning and deep learning algorithms—including Random Forest, XGBoost, CatBoost, Multilater Perceptron Networks, and ensembles—were trained and validated using cross-validation. Systematic band-combination experiments and spatial aggregation strategies were tested to optimize prediction. Results show depth-dependent performance. At the surface, C2X-Complex with XGBoost and ensemble models achieved $R^{2}$ = 0.89; at 1–2 m, CatBoost and ensemble models reached $R^{2}$ = 0.87; at 2–3 m, TOA reflectances with KNN performed best ($R^{2}$ = 0.81); while at 3–4 m, Random Forest achieved $R^{2}$ = 0.66. Generated maps successfully reproduced known eutrophication events (e.g., 2016 crisis, 2025 surge), confirming robustness. The study delivers an end-to-end, validated methodology for depth-specific chlorophyll mapping, surpassing previous surface-only efforts. Its integration of multispectral band combinations, buoy calibration, and ML/DL modeling offers a transferable framework for other turbid coastal systems.
publicationid: pubid: 0000–0000/00\$00.00 © 2025 IEEEfootnotetext: 1Department of Information and Communication Engineering, University of Murcia, Murcia, 30100, Spain.footnotetext: 2Corresponding author. Email: antonio.martinezi@um.esfootnotetext: This work is supported by project NEREIDAS TSI-100120-2024-13 funded by EU NextGenerationEU (PRTR). The study is also supported by grant RYC2023-043553-I , funded by MICIU/AEI/10.13039/501100011033 and ESF++ and by the HORIZON-MSCA-2021-SE-01-01 project Cloudstars (g.a. 101086248).footnotetext: This work has been submitted to the IEEE Transactions on Geoscience and Remote Sensing for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
I Introduction
@@
The Mar Menor, located in southeastern Spain’s Region of Murcia, is Europe’s largest coastal saltwater lagoon. It covers an area of approximately 135 km² and has shallow waters with a maximum depth of seven meters. It has historically been characterized by clear waters, high salinity, and an oligotrophic nature. The lagoon is recognized as a unique and ecologically valuable system and is protected under multiple international designations [1, 2]. However, since the mid-20th century, the lagoon has experienced increasing environmental degradation due to nutrient inputs from intensive agriculture, urban expansion, and former mining activity in the watershed. This chronic pollution has led to the progressive accumulation of nutrients, resulting in the first major eutrophication crisis in 2016 [3]. The resulting phytoplankton bloom, known as the “green soup”, caused widespread turbidity and the collapse of submerged vegetation, marking a turning point in the system’s ecological trajectory. Subsequent episodes of hypoxia and mass mortality of aquatic fauna, particularly in 2019 and 2021, further confirmed the lagoon’s loss of resilience. Even though there was some recovery in chlorophyll levels from 2022 to 2024, a new surge was detected in the summer of 2025 [4]. Concentrations of Chl-a reached 4-5 mg/m³, increasing water turbidity and raising renewed concerns about the risk of anoxia and eutrophication. The recurrence and unpredictability of these eutrophication episodes underscore the need for improved monitoring strategies, as traditional approaches often lack the spatial and temporal resolution required to capture rapid ecological changes. In recent decades, the IoT paradigm has undergone significant expansion, especially in the field of environmental monitoring [5, 6]. The proliferation of connected sensors, capable of collecting data in real time and transmitting it remotely, is transforming the way natural ecosystems are observed and managed. This trend responds to the need for denser, automated and continuous observation systems that allow early detection of environmental changes with high spatiotemporal resolution.
@@
Particularly in vulnerable contexts, the deployment of IoT sensor networks has become a key tool to ensure adaptive, data-driven management. After the 2016 crisis, a network of multiparametric buoys was deployed in the Mar Menor, allowing the measurement of key water quality parameters, such as temperature, salinity, dissolved oxygen, and chlorophyll-a (Chl-a). However, this approach faces operational challenges, such as sensor maintenance and biofouling. In addition, other sensors were deployed in nearby watercourses to measure streamflow and precipitation [7]. Based on those, recent studies have also explored AI-based approaches to predict hydrological dynamics in the Mar Menor basin, with a focus on short-term streamflow forecasting in the Albujón watercourse [8].
@@
Complementing in situ instrumentation, remote sensing has become a valuable tool for the spatial and temporal monitoring of water quality in the Mar Menor, particularly Chl-a. Parameters such as Chl-a concentration (an indicator of phytoplankton biomass) and turbidity have optical properties that allow them to be studied using satellite images. Indeed, algal pigments alter the reflectance of water at certain wavelengths [9], so that multispectral sensors on board satellites can be used to estimate surface Chl-a concentration from reflected radiation. This capability offers notable advantages: unlike buoys, satellite images cover the entire area of the lagoon, allowing the spatial distribution of algal blooms to be mapped and the most affected areas to be detected. In addition, modern satellites have orbits that span a few days, which facilitates more frequent monitoring. Therefore, the use of satellite imagery provides a synergistic monitoring system alongside direct measurements: while buoys provide data at fixed points, remote sensing provides a synchronous view of the entire lagoon. This multi-parametric and multi-scale integration is especially useful in a fragile ecosystem such as the Mar Menor, as it facilitates the early detection of algal blooms and the monitoring of their evolution.
@@
The objective of this study is to create a reliable model of Mar Menor’s Chl-a concentration using satellite imagery to improve spatial and temporal resolution for consistent monitoring. To that end, different sources of data are used as input: in situ data from buoys and satellite images gathered from the European Space Agency [10]. Once a set of reliable, tested models is developed and actively working, Mar Menor monitoring will be more comprehensive, and the output chlorophyll maps will help achieve the goal of making data more accessible to the public.
@@
  • An in-depth exploration of multispectral band combinations as predictors evaluated with a wide array of machine and deep learning models.
  • An end-to-end methodology for predicting Chl-a in turbid waters environments, leveraging the C2RCC processors for atmospheric correction of Sentinel 2 imagery and using buoy data as ground truth to train the models. The methodology has been validated both quantitatively, through performance metrics, and qualitatively, by reproducing known states of the lagoon.
@@
The structure of the paper is as follows: Section II describes the background knowledge needed to contextualize and understand this study and discusses other works that use remote sensing in similar scenarios. Then, in Section III the study area is described, along with the satellite imagery and buoy data. Section IV details the procedure carried out to gather, process and model the data. Section V describes the experiments and results, and finally Section VI highlights the outcomes of this study and proposes future lines of work.
II Background and Related Work
@@
II-A Background
II-A1 Mar Menor
The Mar Menor, located in the Region of Murcia, in southeastern Spain, is the largest coastal saltwater lagoon in Europe, with an area of about 135 km2 and shallow waters, less than 7 meters deep.
@@ II-A1 Mar Menor
Until a few years ago, it was characterized by its crystalline waters, high salinity and oligotrophic nature (very low nutrients disolved), which made it a unique ecosystem of great environmental value. Actually, the ecological importance of the Mar Menor has been recognized with multiple protection figures (regional, state, European and international), including its designation as a Specially Protected Area of Mediterranean Importance (Barcelona Convention) [1], as well as a Site of Community Importance and a Special Protection Area for Birds within the Natura 2000 Network [2].
@@
Despite its natural value, the lagoon has suffered a gradual deterioration over the last decades. Since the mid-20th century, there has been intense socioeconomic development in its watershed with the expansion of irrigated agriculture, intensive livestock farming, urban-tourist growth and former mining operations nearby. These activities have exerted sustained polluting pressures for decades, discharging nutrients such as nitrates and phosphates, among others, into the lagoon through seasonal watercourses and the subway aquifer, leading to a progressive eutrophication of the ecosystem. In fact, over the last 40 years the Mar Menor has suffered chronic pollution from agricultural runoff, which has led to large-scale algal blooms and marked ecological degradation. It should be noted that, despite signs of degradation, until well into the 21st century some environmental indicators still showed low nutrient and chlorophyll values, which led to the lagoon being perceived as relatively oligotrophic. However, nutrients continued to accumulate in the system, making it increasingly vulnerable to ecological collapse [4].
@@
The turning point came in 2016 [11], the year in which the Mar Menor suffered its first major eutrophic crisis. During the spring-summer of that year, the phenomenon known as “green soup” took place: an explosive growth of phytoplankton (microalgae) fed by the excess of nutrients in the water column. The algal proliferation was so dense that it completely clouded the waters and prevented the passage of light beyond  3 m depth, causing the mortality of most of the underwater meadows of phanerogams and macroalgae at the bottom of the lagoon. This event marked an unprecedented ecological collapse in the Mar Menor, showing that the nutrient load had exceeded the system’s capacity for self-regulation. During this crisis, Chl-a concentrations peaked over 30 mg/m3. To put this value into perspective, in 2002 a value higher than 7 mg/m3 was considered a maximum [12]. According to [13], a water body is considered oligotrophic when Chl-a concentrations range between 0 and 2.6 mg/m3, mesotrophic between 2.6 and 7.3 mg/m3, and eutrophic above 7.3 mg/m3. Accordingly, in 2002 the lagoon was close to the eutrophic threshold, whereas in 2016 it exceeded it substantially.
@@
In subsequent years, new critical episodes associated with eutrophication occurred. In autumn 2019, following torrential rains produced by a DANA (an isolated high-altitude atmospheric depression that can generate extreme precipitation events in the Mediterranean region), an episode of anoxia was triggered in the waters causing massive mortality rates on fish and crustaceans. This was due to the high concentration of phytoplankton in water surface, keeping light from reaching the bottom and thus preventing photosynthesis to take place, leading to reduced amounts of dissolved oxygen whcih led to caused a massive mortality of fish and crustaceans. More recently, in August 2021, another acute hypoxia event was recorded in several areas of the lagoon, again linked to an exacerbated phytoplankton bloom. This bloom of 2021, originating around the mouth of the Albujón seasonal watercourse, caused again the suffocation of the aquatic biota and the consequent mortality of thousands of organisms. During 2022, 2023 and 2024 chlorophyll values returned to typical values around 0.5-1.0 mg/m³, but despite various management and mitigation efforts since then, the lagoon remains in a fragile state. Proof of this is the situation currently recorded [14], in the summer of 2025, when a worrying rebound in Chl-a has been detected. Chl-a levels in the water have returned to peaks of 4-5 mg/m³. This increase in Chl-a, observed from July 2025, has been accompanied by an increase in turbidity, once again placing the Mar Menor on alert for the risk of another episode of anoxia if the trend persists.
@@
The severe environmental crises experienced, especially since 2016, have highlighted the need to monitor the conditions of the Mar Menor more closely to anticipate critical episodes [15]. In that year, a network of in-situ sensors was implemented, consisting of multiparametric oceanographic buoys distributed throughout the lagoon. Each buoy is equipped with sensors that measure conductivity, temperature, dissolved oxygen, and Chl-a at different depths, but with limited time resolution. Additionally, several campaigns to take in situ measurements have been carried out, obtaining also values for the aforementioned parameters. This intensive monitoring of the Mar Menor represents a significant advance towards its protection, although it is not enough if the causes of eutrophication in its basin are not addressed. The dependence on buoys also poses a problem: they require frequent on-site maintenance to prevent the sensors from being covered by algae, becoming filled with salt, etc.
II-A2 Sentinel 2
@@
Each satellite carries an MSI instrument (Multispectral Instrument) that captures thirteen spectral bands, from visible to near-infrared and shortwave infrared, detailed in Table I. The mission provides images with spatial resolutions of 10 m (4 bands), 20 m (6 bands), and 60 m (3 bands), with a field of view of approximately 290 km [16].
@@
The Sentinel 2 L1C product provides orthorectified images in Top of Atmosphere (TOA) reflectance, i.e. solar radiation reflected back to the satellite without removing atmospheric effects. These products are georeferenced in the UTM/WGS84 system and are organized in regular 100 km × 100 km mosaics, maintaining the native spatial resolution of each band (10, 20 or 60 meters). They include basic geometric and radiometric corrections, as well as various masks that indicate saturated pixels, areas without data or anomalies in the image. Since they do not incorporate atmospheric correction, the L1C products are especially useful for users who wish to apply proprietary algorithms or perform custom processing from raw data, although these products are not directly comparable between dates without additional processing [17].
The L2A product represents a more advanced and directly usable version, as it includes Bottom of Atmosphere (BOA) type reflectance, i.e. the signal reflected from the land or water surface after removing the effects of the atmosphere using the Sen2Cor correction processor [18]. It is also delivered in the same map geometry, UTM, and with the bands in their native resolution. These products also incorporate a pixel classification layer known as Scene Classification Map (SCL), which identifies areas of cloud, cloud shadow, vegetation, bare ground, water or other types of cover, facilitating the filtering of invalid or contaminated data. L2A products are commonly used in environmental applications such as monitoring water bodies, vegetation or land use dynamics.
TABLE I: Sentinel 2 bands
@@
Satellite images need atmospheric corrections in order to obtain reliable reflectances across all the wavelengths measured. The atmospheric correction performed by the L2A product is not enough to obtain an useful product, since water reflectances make the problem highly complicated. One of the main challenges in the study of inland and coastal water bodies by remote sensing is accurately correcting the reflectance measured by satellite sensors for atmospheric effects. This challenge is exacerbated in waters classified as Case 2, which are rich in suspended particles, dissolved organic matter, and phytoplankton. The optical properties of these waters are highly complex and variable. The C2RCC (Case 2 Regional CoastColor) processor [19, 20] offers a solution consisting of neural networks that are trained using an extensive database of radiative transfer simulations. These networks perform atmospheric correction and estimate the Inherent Optical Properties (IOPs) of water from TOA reflectances.
@@
The C2RCC processor excels in its ability to operate in optically complex environments thanks to a bio-optical model that better captures the diversity of optical conditions in inland and coastal waters. Neural networks are trained to invert TOA reflectances and recover water-leaving reflectance. One of the great advantages of C2RCC over other more empirical or semi-analytical approaches is its ability to process images from multiple sensors (Sentinel 2 MSI, Sentinel-3 OLCI, MODIS, MERIS, etc.) without retraining, making it an operational, scalable and flexible tool for environmental monitoring [21].
@@
To extend the applicability of the C2RCC processor to a wider range of optical conditions in inland and coastal waters, variants of the processor based on neural networks trained with different ranges of IOPs have been developed [22]. The original version, C2RCC-Net, is optimized for medium-turbidity coastal waters. The C2X-Net version considerably expands these training ranges by incorporating extreme cases of suspended matter concentration and absorption by pigments and dissolved organic matter. This was accomplished by supplementing the original database with additional simulations from the CoastColour project, including highly eutrophic water conditions and limiting cases. C2X-Net has demonstrated an enhanced capacity to accurately recover water reflectance in eutrophic lakes, turbid estuaries, and intricate inland lagoons, environments in which more conservative algorithms often falter or produce systematic biases. An intermediate version, C2X-Complex (C2XC), was subsequently introduced to balance the sensitivity of C2RCC with the robustness offered by C2X in extreme conditions. C2XC allows for positive results in moderately turbid waters and more productive bodies without the overfitting or loss of spectral accuracy issues that can occur with C2X in clearer scenarios.
@@
The processor is available as an integrated module in the Sentinel Application Platform (SNAP) software developed by ESA [23], allowing its direct application on Sentinel 2 images with an accessible and user-configurable interface.
II-B Related Work
Since the 2016 eutrophication crisis more options for monitoring water quality have been explored. Once the Sentinel 2 constellation, consisting of 2A and 2B, became operational in 2017, remote sensing techniques became much more viable. Prior to that, [24] studied the use of Landsat and Spot satellites to understand the variability and evolution of Chl-a in the Mar Menor. This study examines the evolution of Chl-a in the Mar Menor using in situ and remote sensing data from multispectral sensors on the Landsat 8 Operational Land Imager (OLI) and SPOT 7 satellites. Normalized indexes were used between bands of the OLI sensor in the visible and near-infrared range of the type:

### v2

@@ Abstract
The Mar Menor, Europe’s largest hypersaline coastal lagoon, located in southeastern Spain, has undergone severe eutrophication crises, with devastating impacts on biodiversity and water quality. Monitoring chlorophyll-a, a proxy for phytoplankton biomass, is essential to anticipate harmful algal blooms and guide mitigation strategies. Traditional in situ measurements, while precise, are spatially and temporally limited. Satellite-based approaches provide a more comprehensive view, enabling scalable and long-term monitoring. This study aims to overcome limitations of chlorophyll monitoring, often restricted to surface estimates or limited temporal coverage, by developing a reliable methodology to predict and map chlorophyll-a concentrations across the water column of the Mar Menor. This work integrates Sentinel 2 imagery with buoy-based ground truth to create models capable of high-resolution, depth-specific monitoring, enhancing early-warning capabilities for eutrophication. Nearly a decade of Sentinel 2 images were atmospherically corrected using C2RCC processors. Buoy data were aggregated by depth (0–1 m, 1–2 m, 2–3 m, 3–4 m). Multiple machine learning algorithms, including CatBoost, XGBoost, Support Vector Machines, and Multilayer Perceptron Networks, were trained and validated using a cross-validation scheme with multi-objective optimization functions. Band-combination experiments and spatial aggregation strategies were tested to optimize prediction. The results show depth-dependent performance. The Root Mean Squared Logarithmic Error (RMSLE) obtained ranges from 0.34 at the surface to 0.39 at 3–4 m, while the $R^{2}$ value was 0.76 at the surface, 0.76 at 1–2 m, 0.70 at 2–3 m, and 0.60 at 3–4 m. Generated maps successfully reproduced known eutrophication events (e.g., 2016 crisis, 2025 surge), confirming robustness. The study delivers an end-to-end, validated methodology chlorophyll mapping. Its integration of multispectral band combinations, buoy calibration, and modeling offers a transferable framework for other turbid coastal systems.
keywords:
[…]
1 Introduction
@@
The Mar Menor, located in southeastern Spain’s Region of Murcia, is Europe’s largest hypersaline coastal lagoon. It covers an area of approximately 135 km² and has shallow waters with a maximum depth of seven meters. It has historically been characterized by clear waters, high salinity, and an oligotrophic nature. The lagoon is recognized as a unique and ecologically valuable system and is protected under multiple international designations [SPAMIsRegionalActivity, EUNISSiteFactsheet]. However, since the mid-20th century, the lagoon has experienced increasing environmental degradation due to nutrient inputs from intensive agriculture, urban expansion, and former mining activity in the watershed. This chronic pollution has led to the progressive accumulation of nutrients, resulting in the first major eutrophication crisis in 2016 [InformeIntegralSobre2017]. The resulting phytoplankton bloom, known as the “green soup", caused widespread turbidity and the collapse of submerged vegetation, marking a turning point in the system’s ecological trajectory. Subsequent episodes of hypoxia and mass mortality of aquatic fauna, particularly in 2019 and 2021, further confirmed the lagoon’s loss of resilience. Even though there was some recovery in chlorophyll levels from 2022 to 2024, a new surge was detected in the summer of 2025 [ColapsoMarMenor]. Chlorophyll, particularly chlorophyll-a (Chl-a), is essential for phytoplankton photosynthesis and is commonly used as an indirect indicator of its biomass, as it modifies the optical properties of water and gives it green tones. Concentrations of Chl-a reached 4-5 mg/m³ in the 2025 event, along with increased water turbidity and raising renewed concerns about the risk of anoxia and eutrophication. The recurrence and unpredictability of these eutrophication episodes underscore the need for improved monitoring strategies, as traditional approaches often lack the spatial and temporal resolution required to capture rapid ecological changes [lopez-andreuMonitoringSystemMar2022].
In recent decades, the IoT paradigm has undergone significant expansion, especially in the field of environmental monitoring [aryaGreenerTomorrowIoTEnabled2023, omkarbhagwankhilariIoTBasedEnvironment2024]. The proliferation of connected sensors, capable of collecting data in real time and transmitting it remotely, is transforming the way natural ecosystems are observed and managed. This trend responds to the need for denser, automated and continuous observation systems that allow early detection of environmental changes with high spatiotemporal resolution.
@@
Particularly in vulnerable contexts, the deployment of IoT sensor networks has become a key tool to ensure adaptive, data-driven management. After the 2016 crisis, a network of multiparametric buoys was deployed in the Mar Menor, allowing the measurement of key water quality parameters, such as temperature, salinity, dissolved oxygen, and Chl-a. However, this approach faces operational challenges, such as sensor maintenance and biofouling. In addition, other sensors were deployed in nearby watercourses to measure streamflow and precipitation [VisorSAIH]. Based on those, recent studies have also explored AI-based approaches to predict hydrological dynamics in the Mar Menor basin, with a focus on short-term streamflow forecasting in the Albujón watercourse [cisterna-garciaArtificialIntelligenceStreamflow2025].
@@
Complementing in situ instrumentation, remote sensing has become a valuable tool for the spatial and temporal monitoring of water quality in the Mar Menor, particularly Chl-a. Parameters such as Chl-a concentration, (an indicator of phytoplankton biomass, and turbidity have optical properties that allow them to be studied using satellite images. Indeed, algal pigments alter the reflectance of water at certain wavelengths [parsonsManualChemicalBiological1984], so that multispectral sensors on board satellites can be used to estimate surface Chl-a concentration from reflected radiation. This capability offers notable advantages: unlike buoys, satellite images cover the entire area of the lagoon, allowing the spatial distribution of algal blooms to be mapped and the most affected areas to be detected. In addition, modern satellites have orbits that span a few days, which facilitates more frequent monitoring. Therefore, the use of satellite imagery provides a synergistic monitoring system alongside direct measurements: while buoys provide data at fixed points, remote sensing provides a synchronous view of the entire lagoon. This multi-parametric and multi-scale integration is especially useful in a fragile ecosystem such as the Mar Menor, as it facilitates the early detection of algal blooms and the monitoring of their evolution.
@@
The objective of this study is to create a reliable model of Mar Menor’s Chl-a concentration using satellite imagery to improve spatial and temporal resolution for consistent monitoring. To that end, different sources of data are used as input: in situ data from buoys and satellite images gathered from the European Space Agency [portaldelaagenciaespacialeuropeaESASpain]. Once a set of reliable, tested models is developed and actively working, Mar Menor monitoring will be more comprehensive, and the output chlorophyll maps will help achieve the goal of making data more accessible to the public.
@@
  • An in-depth exploration of multispectral band combinations as predictors evaluated with a wide array of machine learning (ML) models.
  • An end-to-end methodology for predicting Chl-a in turbid waters environments, leveraging the Case 2 Regional CoastColor (C2RCC) processors for atmospheric correction of Sentinel 2 imagery and using buoy data as ground truth to train the models. The methodology has been validated both quantitatively, through performance metrics, and qualitatively, by reproducing known states of the lagoon.
@@
  • A pipeline for automatic chlorophyll map generation that, starting from a selected date, processes Sentinel 2 SAFE files in SNAP, extracts reflectances for all pixels in the Mar Menor, incorporates additional features, such as band combinations, and performs model inference to generate chlorophyll maps in a fully automated manner. Available in the public Github repository of this project [^1] https://github.com/Antonio-MI/mar-menor-chl.
The structure of the paper is as follows: Section 2 describes the background knowledge needed to contextualize and understand this study and discusses other works that use remote sensing in similar scenarios. Then, in Section 3 the study area is described, along with the satellite imagery and buoy data. Section 4 details the procedure carried out to gather, process and model the data. Section 5 describes the experiments and results, and finally Section 6 highlights the outcomes of this study and proposes future lines of work.
2 Background and Related Work
@@
2.1 Background
2.1.1 Mar Menor
@@
The Mar Menor, located in the Region of Murcia, in southeastern Spain, is the largest hypersaline coastal lagoon in Europe, and until a few years ago, it was characterized by its crystalline waters, high salinity and oligotrophic nature (very low nutrients disolved), which made it a unique ecosystem of great environmental value. Actually, the ecological importance of the Mar Menor has been recognized with multiple protection figures (regional, state, European and international), including its designation as a Specially Protected Area of Mediterranean Importance (Barcelona Convention) [SPAMIsRegionalActivity], as well as a Site of Community Importance and a Special Protection Area for Birds within the Natura 2000 Network [EUNISSiteFactsheet].
@@
Despite its natural value, the lagoon has suffered a gradual deterioration over the last decades. Since the mid-20th century, there has been intense socioeconomic development in its watershed with the expansion of irrigated agriculture, intensive livestock farming, urban-tourist growth and former mining operations nearby. These activities have exerted sustained polluting pressures for decades, discharging nutrients such as nitrates and phosphates, among others, into the lagoon through seasonal watercourses and the subway aquifer, leading to a progressive eutrophication of the ecosystem. In fact, over the last 40 years the Mar Menor has suffered chronic pollution from agricultural runoff, which has led to large-scale algal blooms and marked ecological degradation. It should be noted that, despite signs of degradation, until well into the 21st century some environmental indicators still showed low nutrient and chlorophyll values, which led to the lagoon being perceived as relatively oligotrophic. However, nutrients continued to accumulate in the system, making it increasingly vulnerable to ecological collapse [ColapsoMarMenor].
@@
The turning point came in 2016 [ProblematicaActual], the year in which the Mar Menor suffered its first major eutrophic crisis. During the spring-summer of that year, the phenomenon known as “green soup" took place: an explosive growth of phytoplankton (microalgae) fed by the excess of nutrients in the water column. The algal proliferation was so dense that it completely clouded the waters and prevented the passage of light beyond  3 m depth, causing the mortality of most of the underwater meadows of phanerogams and macroalgae at the bottom of the lagoon. This event marked an unprecedented ecological collapse in the Mar Menor, showing that the nutrient load had exceeded the system’s capacity for self-regulation. During this crisis, Chl-a concentrations peaked over 30 mg/m3. To put this value into perspective, in 2002 a value higher than 7 mg/m3 was considered a maximum [perez-ruzafaSpatialTemporalVariations2005]. According to [sheelaEnvironmentalStatusTropical2011], a water body is considered oligotrophic when Chl-a concentrations range between 0 and 2.6 mg/m3, mesotrophic between 2.6 and 7.3 mg/m3, and eutrophic above 7.3 mg/m3. Accordingly, in 2002 the lagoon was close to the eutrophic threshold, whereas in 2016 it exceeded it substantially.
@@
In subsequent years, new critical episodes associated with eutrophication occurred. In autumn 2019, following torrential rains produced by a DANA (an isolated high-altitude atmospheric depression that can generate extreme precipitation events in the Mediterranean region), an episode of anoxia was triggered in the waters causing massive mortality rates on fish and crustaceans. This was due to the high concentration of phytoplankton in water surface, keeping light from reaching the bottom and thus preventing photosynthesis to take place, leading to reduced amounts of dissolved oxygen whcih led to caused a massive mortality of fish and crustaceans. More recently, in August 2021, another acute hypoxia event was recorded in several areas of the lagoon, again linked to an exacerbated phytoplankton bloom. This bloom of 2021, originating around the mouth of the Albujón seasonal watercourse, caused again the suffocation of the aquatic biota and the consequent mortality of thousands of organisms. During 2022, 2023 and 2024 chlorophyll values returned to typical values around 0.5-1.0 mg/m³, but despite various management and mitigation efforts since then, the lagoon remains in a fragile state. Proof of this is the situation currently recorded [pressMarMenorEntra2025], in the summer of 2025, when a worrying rebound in Chl-a has been detected. Chl-a levels in the water have returned to peaks of 4-5 mg/m³. This increase in Chl-a, observed from July 2025, has been accompanied by an increase in turbidity, once again placing the Mar Menor on alert for the risk of another episode of anoxia if the trend persists.
@@
The severe environmental crises experienced, especially since 2016, have highlighted the need to monitor the conditions of the Mar Menor more closely to anticipate critical episodes [erena2017analisis]. In that year, a network of in-situ sensors was implemented, consisting of multiparametric oceanographic buoys distributed throughout the lagoon. Each buoy is equipped with sensors that measure conductivity, temperature, dissolved oxygen, and Chl-a at different depths, but with limited time resolution. Additionally, several campaigns to take in situ measurements have been carried out, obtaining also values for the aforementioned parameters. This intensive monitoring of the Mar Menor represents a significant advance towards its protection, although it is not enough if the causes of eutrophication in its basin are not addressed. The dependence on buoys also poses a problem: they require frequent on-site maintenance to prevent the sensors from being covered by algae, becoming filled with salt, etc.
2.1.2 Sentinel 2
@@
Each satellite carries an MSI instrument (Multispectral Instrument) that captures thirteen spectral bands, from visible to near-infrared and shortwave infrared, detailed in Table 1. The mission provides images with spatial resolutions of 10 m (4 bands), 20 m (6 bands), and 60 m (3 bands), with a field of view of approximately 290 km [europeanspaceagencyS2Mission].
@@
The Sentinel 2 L1C product provides orthorectified images in Top of Atmosphere (TOA) reflectance, i.e. solar radiation reflected back to the satellite without removing atmospheric effects. These products are georeferenced in the UTM/WGS84 system and are organized in regular 100 km $\times$ 100 km mosaics, maintaining the native spatial resolution of each band (10, 20 or 60 meters). They include basic geometric and radiometric corrections, as well as various masks that indicate saturated pixels, areas without data or anomalies in the image. Since they do not incorporate atmospheric correction, the L1C products are especially useful for users who wish to apply proprietary algorithms or perform custom processing from raw data, although these products are not directly comparable between dates without additional processing [europeanspaceagencyS2Products].
The L2A product represents a more advanced and directly usable version, as it includes Bottom of Atmosphere (BOA) type reflectance, i.e. the signal reflected from the land or water surface after removing the effects of the atmosphere using the Sen2Cor correction processor [sentinelonlineCollection0Level2A]. It is also delivered in the same map geometry, UTM, and with the bands in their native resolution. These products also incorporate a pixel classification layer known as Scene Classification Map (SCL), which identifies areas of cloud, cloud shadow, vegetation, bare ground, water or other types of cover, facilitating the filtering of invalid or contaminated data. L2A products are commonly used in environmental applications such as monitoring water bodies, vegetation or land use dynamics.
Table 1: Sentinel 2 bands
@@
Satellite images need atmospheric corrections in order to obtain reliable reflectances across all the wavelengths measured. The atmospheric correction performed by the L2A product is not enough to obtain an useful product, since water reflectances make the problem highly complicated. One of the main challenges in the study of inland and coastal water bodies by remote sensing is accurately correcting the reflectance measured by satellite sensors for atmospheric effects. This challenge is exacerbated in waters classified as Case 2, which are rich in suspended particles, dissolved organic matter, and phytoplankton. The optical properties of these waters are highly complex and variable. The C2RCC (Case 2 Regional CoastColor) processor [doerfferMERISCase22007, c2rcccommunityprojectDocumentationC2RCC] offers a solution consisting of neural networks that are trained using an extensive database of radiative transfer simulations. These networks perform atmospheric correction and estimate the Inherent Optical Properties (IOPs) of water from TOA reflectances. This is needed because at satellite altitude most of sensor-measured signal comes from the surface and the atmosphere [wangAtmosphericCorrectionOcean2014].
@@
The C2RCC processor excels in its ability to operate in optically complex environments thanks to a bio-optical model that better captures the diversity of optical conditions in inland and coastal waters. Neural networks are trained to invert TOA reflectances and recover water-leaving reflectance. One of the great advantages of C2RCC over other more empirical or semi-analytical approaches is its ability to process images from multiple sensors (Sentinel 2 MSI, Sentinel-3 OLCI, MODIS, MERIS, etc.) without retraining, making it an operational, scalable and flexible tool for environmental monitoring [brockmannEvolutionC2RCCNeural2016].
@@
To extend the applicability of the C2RCC processor to a wider range of optical conditions in inland and coastal waters, variants of the processor based on neural networks trained with different ranges of IOPs have been developed [soriano-gonzalezCombinationC2RCCProcessors2022]. The original version, C2RCC-Net, is optimized for medium-turbidity coastal waters. The C2X-Net version considerably expands these training ranges by incorporating extreme cases of suspended matter concentration and absorption by pigments and dissolved organic matter. This was accomplished by supplementing the original database with additional simulations from the CoastColour project, including highly eutrophic water conditions and limiting cases. C2X-Net has demonstrated an enhanced capacity to accurately recover water reflectance in eutrophic lakes, turbid estuaries, and intricate inland lagoons, environments in which more conservative algorithms often falter or produce systematic biases. An intermediate version, C2X-Complex (C2XC), was subsequently introduced to balance the sensitivity of C2RCC with the robustness offered by C2X in extreme conditions [soriano-gonzalezCombinationC2RCCProcessors2022]. C2XC allows for positive results in moderately turbid waters and more productive bodies without the overfitting or loss of spectral accuracy issues that can occur with C2X in clearer scenarios.
@@
The processor is available as an integrated module in the Sentinel Application Platform (SNAP) software developed by ESA [SNAPEarthOnline], allowing its direct application on Sentinel 2 images with an accessible and user-configurable interface.
2.2 Related Work
Since the 2016 eutrophication crisis more options for monitoring water quality have been explored. Once the Sentinel 2 constellation, consisting of 2A and 2B, became operational in 2017, remote sensing techniques became much more viable. Prior to that, [erenaAnalisisMultisensorVariabilidad2017] studied the use of Landsat and Spot satellites to understand the variability and evolution of Chl-a in the Mar Menor. This study examines the evolution of Chl-a in the Mar Menor using in situ and remote sensing data from multispectral sensors on the Landsat 8 Operational Land Imager (OLI) and SPOT 7 satellites. Normalized indexes were used between bands of the OLI sensor in the visible and near-infrared range of the type: